lim= (X5 + 3X2 + 3) / (99X2 + 44X + 13)

(x→∞)


1
Expert's answer
2021-01-11T17:34:18-0500

"\lim \limits_{x\to \infty} \frac{x^5+3x^2+3}{99x^2+44x+13}=\lim \limits_{x\to \infty} \frac{1+\frac{3}{x^3}+\frac{3}{x^5}}{\frac{99}{x^3}+\frac{44}{x^4}+\frac{13}{x^5}}=\infty"

As we know "\lim \limits_{x\to \infty} \frac{1}{x^n}=0," where "n=1,2,3,4,...." . So expressions in the numerator and denominator in our expression go to zero. But in denominator it goes to zero faster. Therefore, the expression increases indefinitely.